OpenStudy (anonymous):

Duke earned $250 washing cars last summer. He spent it on a stero and a new Jacket. If he paid $190 for a stero what percent of his summer earnings did duke spend on his jacket?

OpenStudy (anonymous):

24% 250-190 = 60/250 = 24%